OpenStudy (anonymous):

The exact value of sin13pi/8 is?

OpenStudy (tkhunny):

Same as \(\sin\left(\dfrac{-3\pi}{8}\right)\)? Same as \(-\sin\left(\dfrac{3\pi}{8}\right)\)? Same as \(-\sin\left(\dfrac{1}{2}\cdot\dfrac{3\pi}{4}\right)\)?

OpenStudy (anonymous):

so would that be -square root of 2-square root of 2/4?

OpenStudy (tkhunny):

No. You will need a half-angle formula.